OpenStudy (anonymous):

what is the greatest common factor of the polynomial below? 7x-21^2

geerky42 (geerky42):

Well, 21 = 7*3

geerky42 (geerky42):

So \(7x - 21^2 = 7x-3^2\cdot 7^2\) Now can you tell me what is GCF here?